OpenStudy (anonymous):

find the width of a rectangle if the diagonal is 18, and the length is 8

OpenStudy (anonymous):

think of the diagonal separating the rectangle into right triangles. you have the hypotenuse (diagonal) and a leg (the length) to find the with use pythagoreon theorem 8^2+b^2=18^2 solve for b

OpenStudy (anonymous):

would be equal b=388

OpenStudy (anonymous):

don't forget to un square root it

OpenStudy (anonymous):

may i get an example how to do this work ?

OpenStudy (anonymous):

64+b^2=324 b^2=260 b= sqrt{260}

OpenStudy (anonymous):

how you get 260 on the example did u divide something

OpenStudy (anonymous):

wait if u subtract that wouldnt u get -260

OpenStudy (anonymous):

324-64=260

OpenStudy (anonymous):

thank u caleb